OpenStudy (anonymous):

Which of the following would be an acceptable first step in simplifying the expression (sinx)/(1 + sinx) ? A. 1/(cscx + sinx) B. sinx + 1 C. (sinx(1 - sinx))/((1 + sinx)(1 - sinx))

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

@zzr0ck3r @terenzreignz

OpenStudy (anonymous):

@Hero @radar @bahrom7893 @thomaster

hero (hero):

Yes, C should be correct.

hero (hero):

The other choices don't really make much sense. They don't represent "first steps"

hero (hero):

B is definitely not a proper first step because you cannot split the denominator in that manner. \[\frac{\sin(x)}{1 + \sin(x)} \ne \frac{\sin(x)}{1} + \frac{\sin(x)}{\sin(x)}\]